# An Efficient Global Optimization Algorithm with Adaptive Estimates of the Local Lipschitz Constants

### v3

@@ 5.3 Hyperparameter Sensitivity Analysis
(a) $\text{HALO}_{\text{L-BFGS-B}}$.
(b) $\text{HALO}_{\text{SDBOX}}$.
(a) $\text{HALO}_{\text{L-BFGS-B}}$.
(b) $\text{HALO}_{\text{SDBOX}}$.
Figure 12: Percentage of problems solved varying the dimensionality $N$ conditional to the value of $\beta$ for the two versions of HALO.
(a) $\text{HALO}_{\text{L-BFGS-B}}$.
(b) $\text{HALO}_{\text{SDBOX}}$.
Figure 13: Average number of functions evaluations varying the dimensionality $N$ conditional to the value of $\beta$ for the two versions of HALO.
5.4 Highlighting the Variable Importance and Problem Interpretability
In addition to its optimization capabilities, HALO offers the potential for extracting valuable insights from the black-box objective function under consideration. In many practical applications, it is not only important to identify the global minimum but also to gain a deeper understanding of the variables that significantly impact the reduction of the objective function. Traditionally, such insights are obtained through sensitivity analysis, which involves computing gradients of the objective function. However, this process can be time-consuming, especially when dealing with computationally expensive simulations.
With HALO, we have an alternative approach to extracting similar information without the need for explicit and additional gradient computations. Once the optimization routine is completed, we can analyze the matrix $\mathbf{G}$ (introduced in Section 4.5), which collects the vectors ~$\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ around each centroid $\mathbf{x}_{i_{k}}$. By examining this matrix, we can gain insights into the directions in which the objective function exhibits greater sensitivity during the iterations of HALO. We can simply determine the variable importance by averaging the rows ~$\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ of the matrix $\mathbf{G}$. This can be expressed as
Variable $\text{Variable Importance}=\frac{1}{|\mathcal{I}_{k}|}\sum_{i}^{|\mathcal{I}_{k}|}\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ (38)
In Fig. 14, we provide an illustrative example of the variable importance for three different objective functions: Eggholder, Dixon Price, and Adjiman respectively in Fig. 14(a), 14(b), and 14(c). By examining the variable importance defined in Eq. 38, we can observe that it accurately identifies the directions in which the objective function exhibits the most significant variation.
(a) Egg Holder.
(b) Dixon Price.
(c) Adjiman.
@@ 5.5 Summary of the Numerical Results and Future Work
  • Highlights problem interpretability: HALO provides valuable insights into the optimization process by identifying the importance of different variables (see Section 5.4), which is crucial for understanding complex objective functions and informing decision-making. This interpretability can also facilitate dimensionality reduction by allowing less significant variables to be excluded from the analysis, simplifying the problem-solving process.
@@
    Computational complexity: We did not compare algorithms based on computational time because, in the context of black-box optimization, the objective function typically originates from a simulation process that constitutes the majority of the computational burden. Function evaluations are often limited, making the computational complexity of the optimization algorithm itself negligible. However, it’s important to note that HALO performs additional operations, such as evaluating local Lipschitz constants, which increases its computational complexity compared to some simpler algorithms. This added complexity can become a limitation when the number of allowed function evaluations is very high (greater than 30,000).
@@
    Sensitive to the problem dimension: Although HALO has demonstrated strong performance across various benchmarks, like any global optimization algorithm, its effectiveness can diminish as the problem dimensionality increases due to the ’curse of the dimensionality’ [61].
@@
  • Considering other partition schemes: Our algorithm uses a central partition scheme to divide the search space but other techniques, such as those proposed in [29], [62] could also be considered.
  • Better convergence properties: The current version of HALO relies on the everywhere dense property, which guarantees to find the global minima as the iterations approach infinity. There may be room for theoretical improvements that offer better convergence properties and more effective stopping criteria. These enhancements should maintain the algorithm’s usability and practicality without introducing crucial hyperparameters that significantly affect convergence.

### v4

@@ 5.5 Summary of the Numerical Results and Future Work
  • Highlights problem interpretability: HALO provides valuable insights into the optimization process by identifying the importance of different variables (see Section 4.5.1), which is crucial for understanding complex objective functions and informing decision-making. This interpretability can also facilitate dimensionality reduction by allowing less significant variables to be excluded from the analysis, simplifying the problem-solving process.
@@
    Computational complexity: We did not compare algorithms based on computational time because, in the context of black-box optimization, the objective function typically originates from a simulation process that constitutes the majority of the computational burden. Function evaluations are often limited, making the computational complexity of the optimization algorithm itself negligible. However, it’s important to note that HALO performs additional operations, such as evaluating local Lipschitz constants, which increases its computational complexity compared to some simpler algorithms. This added complexity can become a limitation when the number of allowed function evaluations is very high (greater than 50,000).
@@
    Sensitive to the problem dimension: Although HALO has demonstrated strong performance across various benchmarks, like any global optimization algorithm, its effectiveness can diminish as the problem dimensionality increases due to the ’curse of the dimensionality’ [bellman1966dynamic]. The current version of HALO is therefore best suited for problems of low to medium dimensionality.
@@
  • Dimensionality Reduction: A promising direction for future research is to use the variable importance derived from HALO to guide dimensionality reduction, which is especially relevant for real-world applications [d2024generative]. One approach would be to leverage the approximated gradients to define an active subspace [constantine2014active, d2024learning], thus focusing on the most influential dimensions of the problem. This can potentially reduce computational costs, make the optimization more efficient, and speed up convergence to the global minimum.
@@
  • Considering other partition schemes: Our algorithm uses a central partition scheme to divide the search space but other techniques, such as those proposed in [paulavivcius2014simplicial], [stripinis2024lipschitz] could also be considered.
  • Better convergence properties: While HALO leverages the everywhere dense property to guarantee a path to the global minimum, future refinements may further strengthen its convergence guarantees or improve stopping criteria. Such improvements should preserve the algorithm’s practical usability and avoid introducing new, critical hyperparameters..
@@ 6 Conclusions
(a) First benchmark.
(b) Second Benchmark.
(c) Third benchmark.
Figure 9: Operational characteristics for the three benchmarks.
(a) First benchmark.
(b) Second benchmark.
(c) Third benchmark.
Figure 10: Percentage of the problem solved varying the dimensionality NN for the three benchmarks.
(a) First benchmark.
(b) Second benchmark.
(c) Third benchmark.
Figure 11: Average number of functions evaluations varying the dimensionality NN for the three benchmarks.
(a) Hard functions.
(b) Simple functions.
(a) $\text{HALO}_{\text{L-BFGS-B}}$.
(b) $\text{HALO}_{\text{SDBOX}}$.
(a) $\text{HALO}_{\text{L-BFGS-B}}$.
(b) $\text{HALO}_{\text{SDBOX}}$.
Figure 14: Percentage of problems solved varying the dimensionality NN conditional to the value of $\beta$ for the two versions of HALO.
(a) $\text{HALO}_{\text{L-BFGS-B}}$.
(b) $\text{HALO}_{\text{SDBOX}}$.
Figure 15: Average number of functions evaluations varying the dimensionality NN conditional to the value of $\beta$ for the two versions of HALO.
@@
Acknowledgment
The author would like to thank Prof. Stefano Lucidi for his valuable advice and guidance throughout this work. Appreciation is also extended to the three anonymous reviewers for their constructive comments, which helped improve the quality of the paper.